Question 8 A firm with a DFL of 2.3 expects EBIT to increase by 5.74%. If our last EPS was $3.57, what will our new EPS be?

Question 9 If a firm has a DCL of 2.73 and expects sales to increase by 6%, how much will EPS increase by?
Answer #1

8.

DFL(Degree of Financial leverage) = 2.3

EBIT increase by 5.74%

Last EPS = $3.57

DFL = \frac{\textup{\% change in EPS}}{\textup{\% change in EBIT}}

\textup{\% change in EPS} = \textup{\% change in EBIT}*DFL

\textup{\% change in EPS} = 0.0574*2.3

\large \textup{\% change in EPS} = 0.132020

Thus, New EPS would be:

\textup{New EPS} = 3.57*(1+\textup{\% change in EPS})

\large \textup{New EPS} = \$4.04

9.

DCL (Degree of combined leverage) = 2.73

% change in sales = 6%

DCL = \frac{\textup{\% change in EPS}}{\textup{\% change in Sales}}

\textup{\% change in EPS} = \textup{\% change in Sales}*DCL

Thus, EPS increase by:

\textup{\% change in EPS} = 0.06*2.73

\large \textup{\% change in EPS} = 16.38\%

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.